# Disentangle and Regularize: Sign Language Production with Articulator-Based Disentanglement and Channel-Aware Regularization

### v2

@@ 5.2 Test Results
Table 5: Performance comparison of the models on the PHOENIX14T TEST Set. †  denotes that the scores are obtained using the back translation model shared in SLRTP-Sign-Production-Evaluation [31].
Models Gloss DTW-MJE↓ BLEU-1↑ BLEU-2↑ BLEU-3↑ BLEU-4↑ ROUGE↑ chrF↑
GT - 0.000 34.43 22.08 16.13 12.81 35.22 34.62
Autoregressive Approaches
PT Base [30] 0.197 6.27 3.33 2.14 1.59 14.52 9.50
@@
NSLP-G w/o Finetuning [12] 0.646 - - - 4.41 - -
NAT-AT [8] 0.177 14.26 9.93 7.11 5.53 18.72 -
NAT-EA [8] 0.146 15.12 10.45 7.99 6.66 19.43
NSVQ+Non-Autoreg. Decoding [29] 0.105 27.74 16.36 11.75 9.20 27.93 -
Ours-DARSLP 0.043 28.91 16.45 11.52 8.86 28.50 28 .84
Ours-DARSLP-KL 0.039 32.33 19.84 14.07 10.90 31.80 32.00
We evaluate the best model configuration on the PHOENIX14T and CSL-Daily test sets. Table 5 presents PHOENIX14T results alongside prior work. The DARSLP model, already outperforms several baselines, demonstrating the effectiveness of using ground-truth encodings from the autoencoder. Adding KL divergence in the second phase significantly improves generalization by reducing overfitting to dominant high-variance patterns (e.g., body motion) and encouraging smoother outputs aligned with latent region distributions. This yields measurable gains: BLEU-4 improves from 8.86 to 10.90 and ROUGE from 28.50 to 31.80, confirming the benefit of aligning predicted embeddings with learned priors.
Sample pose sequences are illustrated in Figure 5. As observed in these qualitative examples, the predicted sequences approximate the overall structure of the ground truth, reflecting key manual and non-manual articulations. While the outputs are not frame-perfect reproductions, they generally preserve the semantic intent and flow of the original signing. Despite relying on a non-autoregressive decoder, the model is able to produce smooth transitions and plausible motion patterns, suggesting that the proposed representation and training approach effectively support expressive sign generation. Importantly, these results are achieved without any gloss supervision or pretrained models, using only the PHOENIX14T training set. This demonstrates the potential of our lightweight, data-efficient framework in producing competitive results under limited supervision.
@@
Figure 5: Pose sequences generated from input sentences. Top: ”am mittwoch zieht von der nordsee regen heran der sich am donnerstag ausbreitet”. Bottom: ”und nun die wettervorhersage für morgen donnerstag den siebzehnten dezember”.
Table 6: Evaluation results on the CSL-Daily TEST Set using the finalized configuration.
Model DTW-MJE↓ BLEU-1↑ BLEU-2↑ BLEU-3↑ BLEU-4↑ ROUGE↑ chrF↑
GT 0.000 21.63 10.90 6.31 4.02 21.81 5.26
DARSLP 0.177 19.13 8.33 4.46 2.73 18.72 4.24
DARSLP-KL 0.176 20.03 9.16 5.25 3.35 19.41 4.61
@@
6 Conclusions
In this work, we presented a gloss-free sign language production framework that integrates context-aware text encoding, a semantically structured latent pose space, and dynamic temporal control within a unified architecture. At the core of our approach is a disentangled representation that separates pose features into meaningful regions, such as hands, face, and body, allowing the model to capture manual and non-manual dynamics more effectively. This design enhances representation quality and supports more interpretable modeling of sign language motion.
Our non-autoregressive transformer architecture leverages sentence-level BERT embeddings alongside a dedicated length prediction module, which dynamically adjusts the output sequence length based on the linguistic complexity of the input. This design facilitates temporal alignment between the source text and the generated sign pose sequences.
Furthermore, we investigate the use of KL divergence to align predicted latent distributions with ground truth embedding statistics, serving as an effective regularizer within the structurally factorized latent pose space. Through two-stage training, we empirically show that introducing KL divergence after initial training helps preserve expressiveness in the learned latent representations and plays a key role in reducing the regression-to-the-mean problem, leading to more diverse and expressive sign motion generation.
Despite the absence of gloss supervision and the use of pretrained foundation models, our framework offers a lightweight yet effective solution. Compared to more complex approaches in the literature, it achieves competitive performance through a compact architecture and a targeted training strategy.
Acknowledgements
This research is funded by TÜBİTAK under the 1001 program (grant no: 124E618). We also acknowledge the EuroHPC Joint Undertaking for awarding us access to the Vega supercomputer at IZUM, Slovenia.

### v3

@@ 4.5 Test Results
Table 4 reports test set results on PHOENIX14T, comparing our method with recent autoregressive and non-autoregressive approaches. The DARSLP baseline already outperforms most prior work without requiring gloss annotations, demonstrating the benefit of disentangled latent encodings. Adding KL divergence during the second training phase further improves generalization, boosting BLEU-4 from 8.72 to 11.07 and ROUGE from 23.25 to 32.55. This confirms that aligning predicted embeddings with learned priors reduces overfitting to dominant high-variance features (e.g., body motion) and yields smoother, more accurate sign pose sequences.
Model G DTW↓ B1↑ B2↑ B3↑ B4↑ chrF↑ ROUGE↑
GT - 0.000 34.43 22.08 16.13 12.81 34.62 35.22
Autoregressive
PT Base [30] 0.197 6.27 3.33 2.14 1.59 9.50 14.52
@@
NSLP-G w/o FT [12] 0.646 - - - 4.41 - -
NAT-AT [8] 0.177 14.26 9.93 7.11 5.53 21.87 18.72
NAT-EA [8] 0.146 15.12 10.45 7.99 6.66 22.98 19.43
NSVQ + Non-AR [29] 0.105 27.74 16.36 11.75 9.20 - 27.93
DARSLP (Ours) 0.0439 24.97 15.91 11.35 8.72 29.67 23.25
DARSLP + KL (Ours) 0.0394 33.17 20.47 14.38 11.07 31.69 32.55
Table 4: PHOENIX14T TEST set comparison. Scores via back translation model [31].
Qualitative examples in Figure 7 show that the model captures key manual and non-manual articulations. While the outputs are not frame-perfect reproductions, they generally preserve the semantic intent and flow of the original signing. Despite relying on a non-autoregressive decoder, the model is able to produce smooth transitions and plausible motion patterns, suggesting that the proposed representation and training approach effectively support expressive sign generation. These results, achieved without gloss supervision or large pretrained models, demonstrate the effectiveness and efficiency of our proposed framework under limited supervision.
@@
Figure 7: Pose sequences generated from input sentences. Top: ”am mittwoch zieht von der nordsee regen heran der sich am donnerstag ausbreitet”. Bottom: ”und nun die wettervorhersage für morgen freitag den vierundzwanzigsten juli”.
@@
Model DTW-MJE↓ B1↑ B2↑ B3↑ B4↑ chrF↑ R↑
GT 0.000 21.81 10.90 6.31 4.02 5.26 21.63
DARSLP 0.177 18.72 8.33 4.46 2.73 4.24 19.13
DARSLP+KL 0.175 20.47 9.55 5.45 3.53 4.89 20.55
Table 5: Evaluation results on the CSL-Daily TEST set.
5 Conclusions and Future Work
We proposed a gloss-free sign language production framework that combines context-aware text encoding, a semantically structured latent pose space, and dynamic temporal control within a unified architecture. Despite not relying on gloss annotations or large pretrained generative models, the proposed approach achieves competitive performance through a compact design and targeted training strategy.
Beyond its practical performance, this framework offers a promising research direction by enabling interpretable analysis of region-wise latent behavior. Our visualizations demonstrate how the model responds to variance in specific articulators, offering key insights into the internal dynamics of sign language generation. To the best of our knowledge, this is the first work to present such a targeted, interpretable latent space analysis in the context of sign language production. As such, it also contributes to the broader goal of interpretable AI in sequence generation models.
While qualitative results demonstrate smooth and plausible motions, there is still room to improve visual fidelity and expressiveness. Current evaluation of SLP models relies heavily on back-translation, which introduces additional uncertainty. More direct and reliable evaluation techniques are needed to assess generation quality more faithfully. In parallel with improving evaluation, further architectural refinements and training strategies are also needed. Although this study adopts a region-wise latent structure and disentangled modeling across articulators, assigning appropriate loss weights remains challenging, as each region exhibits different levels of motion complexity and semantic importance. Future work will explore adaptive weighting schemes that better reflect joint-level contributions, particularly to capture the hierarchical motion structure of hands more effectively.